# Aligning Text, Images, and 3D Structure Token-by-Token

### v1

@@ Abstract
Creating machines capable of understanding the world in 3D is essential in assisting designers that build and edit 3D environments and robots navigating and interacting within a three-dimensional space. Inspired by advances in language and image modeling, we investigate the potential of autoregressive models for a new modality: structured 3D scenes. To this end, we propose a unified LLM framework that aligns language, images, and 3D scenes and provide a detailed “cookbook” outlining critical design choices for achieving optimal training and performance addressing key questions related to data representation, modality-specific objectives, and more. We evaluate performance across four core 3D tasks – rendering, recognition, instruction-following, and question-answering – and four 3D datasets, synthetic and real-world. We extend our approach to reconstruct complex 3D object shapes by enriching our 3D modality with quantized shape encodings, and show our model’s effectiveness on real-world 3D object recognition tasks. Project webpage: https://glab-caltech.github.io/kyvo/.
@@ 1 Introduction
Large language models (LLMs) that fuse text and images have unlocked unprecedented visual capabilities, such as visual captioning and text-guided image generation. Inspired by this success, we explore a third modality – structured 3D scenes – and show how aligning it with text and images allows LLMs to tackle a new suite of visual tasks in 3D, such as 3D reconstruction from a single image and 3D-conditioned image generation.
@@
But, how does one design and train an LLM that aligns this structured 3D modality with image and text? While there is extensive work on language-only achiam2023gpt ; grattafiori2024llama3herdmodels ; team2023gemini or vision-and-language models tong2024cambrian ; laurenccon2024matters ; beyer2024paligemma , there is limited work on training with an additional structured 3D modality. We explore the vast design space and evaluate the impact of design choices in architecture, training strategies, data representation, modality-specific objectives, and more. Our testbed consists of four core 3D tasks – image generation, recognition, instruction-following, and question-answering – and four challenging datasets, both synthetic and real-world. Through an extensive “cookbook” of rigorously validated empirical findings we hope to provide guidelines to design 3D-aligned multi-modal LLMs. We then extend our findings to demonstrate the effectiveness of our unified LLM framework, Kyvo (Greek for “3D cube"), to single-view 3D reconstruction of complex object geometries via a tokenization technique for 3D shape encodings. Finally, we apply our model to real-world domains of indoor and outdoor scenes for the task of image-based 3D object recognition, where we show that our model can outperform prior vision specialists designed for the task.
@@ 2 Related Work
LLMs. The success of LLMs is attributed to the scalability of transformers vaswani2017attention taking the form of encoder-decoder tay2022ul2 ; chung2024scaling and most popular decoder-only achiam2023gpt ; grattafiori2024llama3herdmodels ; team2023gemini models. LLMs showcase the effectiveness of autoregressive formulations in achieving task generalization, a finding we extend to 3D structured representations.
Vision-Language Models (VLMs). Early VLMs radford2021learning ; jia2021scaling ; alayrac2022flamingo pioneered learning from image-text pairs, focusing on zero-shot classification, cross-modal retrieval, and question-answering. With the advent of LLMs, modern VLMs liu2024visual ; chen2024internvl ; lu2024deepseek ; ma2024janusflow ; wang2024qwen2 ; wu2024deepseek expand these capabilities, leveraging internet-scale training data and improved alignment strategies. Several works laurenccon2024matters ; beyer2024paligemma ; tong2024cambrian conduct vision-centric analysis and investigate key design choices for training VLMs. However, they limit images only as input. Others team2024chameleon ; zhou2024transfusion ; tong2024metamorph explore early fusion and diffusion for images along with text to support image generation.
LLMs & 3D. Prior work integrates LLMs with various 3D formats for tasks like VQA and semantic grounding ma2024llms . Some ma2024spatialpin ; cheng2024spatialrgpt ; liao2024reasoning extend VLMs to 3D reasoning via depth-aware image queries. Others hong20233d ; tang2024minigpt ; xu2025pointllm ; ma2024find ; marsili2025visual ; kerr2023lerf ; shen2023distilled embed 3D data (e.g., point clouds, NeRFs) with CLIP-style features to support free-form QA. Specifically, 3D-LLM hong20233d inputs holistic 3D scene point clouds (obtained from scanning techniques or structure-from-motion) and trains a language model for the task of language captioning and QA. In contrast, our model, via its structured 3D modality, decomposes scenes into 3D objects, enabling direct alignment with language and vision. This allows broader tasks like 3D shape and pose prediction from a single image and image generation from object-centric 3D scene input. SceneScript avetisyan2025scenescript uses an autoregressive decoder over videos and holistic scene point clouds to predict 3D bounding boxes. In contrast, our method focuses on aligning images with structured 3D scene and object representations.
@@ 3 Building our model token-by-token
We present our approach, Kyvo, that aligns the 3D modality with text and images. The result is a unified multimodal LLM framework than can perform a range of visual 3D tasks, as shown in Fig. 2.
Sec. 3.1 outlines our experimental setup, including tasks and datasets. Sec. 3.2 presents a bottom-up analysis of key design choices – covering data representation, sequence design, and scaling – and addresses challenges in modality-specific tokenization and optimization. We summarize empirical insights and adopt optimal configurations for each subsequent experiment. Finally, we explore generalization to complex shapes (Sec. 3.3), 3D shape representations (Sec. 3.4), and real-world recognition (Sec. 3.5).
@@ 3.1 Setup: tasks and datasets
We design and train an autoregressive model aligning 3D with images and language, capable of performing complex 3D tasks.
Tasks. We focus on four core 3D tasks: image generation from 3D scene specifications (rendering), 3D object recognition from a single image, instruction-following 3D editing and question-answering.
Rendering: 3D $\rightarrow$ Image. Given a 3D scene described by our structured modality (object types, shapes, locations, poses), the model generates a corresponding image. This task tests whether rendering – typically requiring tools (e.g., Blender) and complex processes (e.g., ray tracing, rasterization) – can be reframed as feed-forward next-token prediction.
Recognition: Image $\rightarrow$ 3D. This is a dual of rendering. Given an image, the model predicts the underlying 3D scene structure, including object shapes, types, positions and poses.
[…]
All four tasks require spatial reasoning and grounding. Current VLMs struggle with 3D and geometry prediction, and both rendering and recognition from a single image remain challenging due to issues like object misplacement and hallucinations. Failure cases are shown in Appendix F.
Datasets. We consider four datasets for our experiments: CLEVR johnson2017clevr features scenes of simple shapes in varying layouts; ObjaWorld features complex objects of any geometry; Objectron ahmadyan2021objectron and ARKitScenes dehghan2021arkitscenes comprising real-world indoor and outdoor scenes of various object types.
Our choice of datasets reflects varying complexity and practical use cases. We use CLEVR and design ObjaWorld to simulate the workflow of a design studio that composes and edits 3D scenes with 3D assets. Our autoregressive model provides an alternative to specialized 3D software by enabling image generation, scene inference, and editing in a single forward pass of an LLM, leveraging our structured 3D modality. ObjaWorld showcases our model’s ability to generalize to complex shapes, supporting both asset-level geometry prediction and full scene reconstruction. Finally, Objectron and ARKitScenes demonstrate real-world applicability in robotics, AR, and assistive technologies.
@@
Recognition. Here, the model outputs 3D scenes, evaluated using the Jaccard Index ($J=\frac{tp}{tp+fp+fn}$), which compares predicted and ground-truth objects based on matching attributes (type, size, color, material) and spatial proximity within a threshold $\tau$. True positives (tp) are matched objects within $\tau$; false positives (fp) are unmatched predictions; and false negatives (fn) are missed ground-truth objects. This metric captures both object identification and spatial accuracy, penalizing omissions and hallucinations. We report the average Jaccard Index over $\tau\in\{0.05,0.10,0.15,0.20,0.25\}$.
Question-Answering. We report answer accuracy based on exact match between predicted and ground-truth text (typically 1–2 tokens). A random baseline yields $0.359$ accuracy, while a frequency-based baseline reaches $0.430$. See Appendix C.3 for more details.
Rendering. In the rendering task, images are predicted from 3D scene inputs. Standard image metrics (L2, SSIMwang2004image , FID heusel2017gans , PSNR) fall short, as they don’t capture object placement and attribute accuracy in the generated images. Instead, we rely on human evaluations, where annotators rank model outputs (anonymized) against the ground-truth scene, and we report the mean rank. While L2 and SSIM trends align with human judgment, we report human evaluations as they directly assess scene correctness. More details are provided in Appendix C.2.
@@ 3.2 Cookbook
Our model, Kyvo, builds on the decoder-only Llama-3.2-1B-Instruct grattafiori2024llama3herdmodels , originally designed for text-only tasks, and extends it with modality-specific tokenizers for images and the structured 3D modality, along with modified input embeddings and output projections (Fig. 1). We evaluate key architectural choices – each with significant performance impact – which we hope offer practical insights to guide future multimodal LLM development in the 3D domain.
We begin by exploring how to represent and tokenize our structured 3D modality using CLEVR, which contains 3–10 object scenes with varied layouts. Each scene is described by object type, color, material, size, and 3D position, and rendered in Blender to generate ground truth for all four tasks – totaling $120,\!000$ scenes. This setup helps establish best practices for integrating 3D into a unified LLM framework. We then extend these findings to complex geometries and shape encodings in ObjaWorld (Sec. 3.3 & Sec. 3.4) and validate generalization to real-world scenes in Objectron and ARKitScenes (Sec. 3.5).
3.2.1 What is the optimal representation for modalities?
Our model handles three modalities – images, text, and structured 3D scenes – by converting each into token sequences for autoregressive modeling. Below, we outline tokenization strategies for each and evaluate optimal design choices.
Text. We employ an off-the-shelf text tokenizer from Llama-3.2 grattafiori2024llama3herdmodels with a $128,\!000$ size vocabulary. Text instructions, questions, and answers are tokenized and enclosed within special, learnable tokens [TEXT-START] and [TEXT-END].
Images. Our framework addresses tasks involving images as both inputs and outputs. To enable image generation within an autoregressive paradigm, we adopt discrete image representations using VQGAN esser2021taming . This approach maps continuous image features to discrete tokens through a learned codebook. Specifically, we fine-tune a pre-trained VQGAN model to optimize the codebook for our visual domain. Image tokens are enclosed within special, learnable tokens [IMAGE-START] and [IMAGE-END] to form a sequence.
3D scene tokenization. We parse 3D scenes, containing 3D object information in the form of structured lists, to construct sequence of tokens. We encode object attributes – size, color, material, shape, and location – by introducing special, learnable tokens: [SIZE], [COLOR], [MATERIAL], [SHAPE], and [LOCATION]. Advanced attributes such as shape representations and pose are introduced later to address more complex tasks. Each object sequence is enclosed by [OBJECT-START] and [OBJECT-END], and the full scene sequence is wrapped by [SCENE-START] and [SCENE-END]. An example scene with two objects is shown below.
@@ 3.2.1 What is the optimal representation for modalities?
The 3D location and orientation of an object are key attributes for understanding 3D spatial relationships and performing grounded 3D actions based on instructions. Thus, accurately encoding coordinates is essential. However, LLMs are inefficient at handling numbers mirzadeh2024gsm ; singh2024tokenization . To mitigate this, we encode each object’s $x,y,z$ coordinates separately as individual tokens, allowing the model to learn distinct embeddings for each coordinate. We discretize the coordinate values using equally spaced bins based on a chosen granularity. We find this granularity to be decisive for performance – if it is too coarse, spatial inaccuracies arise; if it is too fine, it leads to an exponential increase in tokens, with fewer training samples per bin and thus difficulty in learning. Table 1 shows the effect of granularity across the four tasks on CLEVR. We find that a granularity of $0.05$ outperforms the coarser $0.5$ and the finer $0.005$. Fig. 3 compares image generation for the rendering task on the test set across varying levels of granularity.
@@
Furthermore, our tokenization approach substantially improves efficiency by reducing sequence length compared to standard text tokenizers, which typically fragment floating-point values (e.g., "$0.000$" becomes "00", ".", "$000$"). We achieve a mean sequence length reduction from $271.4$ tokens using the standard Llama-3.2 tokenizer to $93.2$ tokens – a 2.91×2.91$\times$2.91 × compression ratio. These efficiency gains directly translate to decreased memory requirements and faster training and inference times. The final vocabulary comprises $129,\!415$ tokens covering all modalities.

### v2

# Aligning Text, Images and 3D Structure Token-by-Token

@@ Abstract
Creating machines capable of understanding the world in 3D is essential in assisting designers that build and edit 3D environments and robots navigating and interacting within a three-dimensional space. Inspired by advances in language and image modeling, we investigate the potential of autoregressive models for a new modality: structured 3D scenes. To this end, we propose a unified LLM framework that aligns language, images, and 3D scenes and provide a detailed “cookbook” outlining critical design choices for achieving optimal training and performance addressing key questions related to data representation, modality-specific objectives, and more. We show how to tokenize complex 3D objects to incorporate into our structured 3D scene modality. We evaluate performance across four core 3D tasks – rendering, recognition, instruction-following, and question-answering – and four 3D datasets, synthetic and real-world. We show our model’s effectiveness on reconstructing complete 3D scenes consisting of complex objects from a single image and on real-world 3D object recognition tasks. Project webpage: https://glab-caltech.github.io/kyvo.
@@ 1 Introduction
Large language models (LLMs) that fuse text and images have unlocked unprecedented visual capabilities, such as visual captioning and text-guided image generation. Inspired by this success, we explore a third modality – structured 3D scenes – and show how aligning it with text and images allows LLMs to tackle a new suite of visual tasks in 3D, such as 3D reconstruction from a single image and 3D-conditioned image generation. We start from a language-pretrained transformer and extend it with a structured 3D modality that is designed to try to leverage its linguistic reasoning and generalization capabilities.
@@
But, how does one design and train an LLM that aligns this structured 3D modality with image and text? While there is extensive work on language-only [achiam2023gpt, grattafiori2024llama3herdmodels, team2023gemini] or vision-and-language models [tong2024cambrian, laurenccon2024matters, beyer2024paligemma], there is limited work on training with an additional structured 3D modality. We explore the vast design space and evaluate the impact of design choices in architecture, training strategies, data representation, modality-specific objectives, and more. Our testbed consists of four core 3D tasks – image generation, recognition, instruction-following, and question-answering – and four challenging datasets, both synthetic and real-world. Through an extensive “cookbook” of rigorously validated empirical findings we hope to provide guidelines to design 3D-aligned multi-modal LLMs, including guidelines on tokenizing complex 3D shapes. We demonstrate the effectiveness of our unified LLM framework, Kyvo (Greek for “3D cube”), on single-view 3D reconstruction of complex object geometries. Finally, we apply our model to real-world domains of indoor and outdoor scenes for the task of image-based 3D object recognition, where we show that our model can tackle tasks that previously required task-specific vision specialists.
@@ 2 Related Work
LLMs. The success of LLMs is attributed to the scalability of transformers [vaswani2017attention] taking the form of encoder-decoder [tay2022ul2, chung2024scaling] and most popular decoder-only [achiam2023gpt, grattafiori2024llama3herdmodels, team2023gemini] models. LLMs showcase the effectiveness of autoregressive formulations in achieving task generalization, a finding we extend to 3D structured representations.
Vision-Language Models (VLMs). Early VLMs [radford2021learning, jia2021scaling, alayrac2022flamingo] learned from image–text pairs for tasks like zero-shot classification, retrieval, and QA. Modern VLMs [liu2024visual, chen2024internvl, lu2024deepseek, ma2024janusflow, wang2024qwen2, wu2024deepseek] extend these abilities using internet-scale data and stronger alignment. Recent studies [laurenccon2024matters, beyer2024paligemma, tong2024cambrian] analyze vision-centric design choices but treat images only as input, while others [team2024chameleon, zhou2024transfusion, tong2024metamorph] explore early fusion and diffusion for image generation. Nonetheless, current VLMs remain weak in 3D reasoning and geometry prediction, tasks that we tackle in this work.
LLMs & 3D. Prior work integrates LLMs with diverse 3D formats for tasks like VQA and grounding [ma2024llms]. Some [ma2024spatialpin, cheng2024spatialrgpt, liao2024reasoning] extend VLMs to 3D via depth-aware queries, while others [hong20233d, tang2024minigpt, xu2025pointllm, ma2024find, marsili2025visual, kerr2023lerf, shen2023distilled] embed 3D data (e.g., point clouds, NeRFs) with CLIP-style features for open-ended QA. 3D-LLM [hong20233d] uses holistic 3D point clouds for captioning and QA, whereas our model’s structured 3D modality decomposes scenes into objects, enabling direct alignment across language and vision. This supports tasks like 3D shape and pose prediction from a single image and image generation from object-centric 3D inputs. Similarly, while SceneScript [avetisyan2025scenescript] autoregressively predicts 3D boxes from videos and full scene clouds, our approach aligns images with structured 3D object and scene representations.
3D Tokenization. Prior work on 3D tokenization targets limited tasks, mainly single-asset generation. SAR3D [chen2024sar3d] uses triplanes, while AToken [lu2025atokenunifiedtokenizervision] builds on Trellis [xiang2024structured] to train a joint tokenizer for images, 3D, and video. In contrast, we design a 3D tokenizer for structured scene encoding, optimized for compactness to represent multiple objects per scene. Whereas AToken and SAR3D use over $20\mathrm{k}$ and $2040$ tokens per asset respectively, our tokenizer achieves comparable or better reconstruction with far fewer tokens (see Sec.˜3.2).
@@ 3 Building our model token-by-token
We present our approach, Kyvo, that aligns the 3D modality with text and images. The result is a unified multimodal LLM framework that can perform a range of visual 3D tasks – rendering, reconstruction, recognition and instruction-following – as shown in Fig.˜2.
Sec.˜3.1 outlines our experimental setup, including tasks and datasets. Sec.˜3.2 presents a bottom-up analysis of key design choices – covering data representation, 3D tokenization, sequence design, and scaling – and addresses challenges in modality-specific tokenization and optimization. We summarize empirical insights and adopt optimal configurations for each subsequent experiment. Finally, we explore generalization to complex scene layouts (Sec.˜3.3), 3D shape representations (Sec.˜3.4), and real-world recognition (Sec.˜3.5).
@@ 3.1 Setup: tasks and datasets
We design and train an autoregressive model aligning 3D with images and language, capable of performing 3D tasks.
Tasks. We focus on four core 3D tasks: image generation from 3D scene specifications (rendering), 3D object recognition & reconstruction from a single image, instruction-following 3D editing and question-answering.
Rendering: 3D → Image. Given a 3D scene described by our structured modality (object types, shapes, locations, poses), the model generates a corresponding image. This task tests whether rendering – typically requiring tools (e.g., Blender) and complex processes (e.g., ray tracing, rasterization) – can be reframed as feed-forward next-token prediction.
Reconstruction: Image → 3D. This is a dual of rendering. Given an image, the model predicts the underlying 3D scene structure, including object shapes, types, positions and poses.
Instruction-Following: $(\text{Image},\text{3D},\text{Text}_{\text{I}})$$(\text{Image},\text{3D})$. We define a set of tasks to manipulate and modify 3D scenes given a text instruction. These include: (1) modifying the appearance of objects, (2) adding new objects, (3) removing objects, and (4) moving an object to a desired location. For each subtask, we craft templated natural language instructions to evaluate the model’s ability to follow instructions in 3D, e.g. “Remove the red mug behind the yellow bottle”.
Question-Answering: $(\text{Image},\text{3D},\text{Text}_{\text{Q}})$$\text{Text}_{\text{A}}$. We generate QA pairs using templated questions. Given the 3D scene, an image, and a query (e.g. “Is there a green object the same size as the glass vase?”), the model predicts a natural language answer. Details are in the Appendix.
All four tasks require spatial reasoning and grounding. Two of them explicitly involve text as input or output and all four implicitly require linguistic reasoning in the use of the structured 3D modality as input or output.
Datasets. We consider four datasets for our experiments. Two synthetic ones: CLEVR [johnson2017clevr] features scenes of simple shapes in varying layouts; ObjaWorld features complex objects of any geometry sourced from Objaverse [deitke2023objaverse]. Two real-world ones: Objectron [ahmadyan2021objectron] and ARKitScenes [dehghan2021arkitscenes] comprising real-world indoor and outdoor scenes of various object types.
@@
Recognition. We evaluate predicted 3D scenes using the Jaccard Index, $J=\frac{tp}{tp+fp+fn}$, which measures object-level agreement between prediction and ground truth based on matching attributes (type, size, color, material) and spatial proximity within threshold $\tau$. True positives are matched objects within $\tau$, false positives are extra predictions, and false negatives are misses. We report the mean Jaccard Index over $\tau\in\{0.05,0.10,0.15,0.20,0.25\}$, reflecting both recognition and spatial accuracy.
Question-Answering. We report answer accuracy based on exact match between predicted and ground-truth text (typically 1–2 tokens). A random baseline yields $0.359$ accuracy, while a frequency-based baseline reaches $0.430$. See Appendix for more details.
Rendering. In the rendering task, models generate images from 3D scene inputs. Standard image metrics (L2, SSIM [wang2004image], FID [heusel2017gans], PSNR) fail to capture object placement or attribute accuracy. We therefore use human evaluations: annotators rank anonymized outputs against ground truth, and we report the mean rank. While L2 and SSIM roughly follow these trends, human judgment better reflects scene correctness (details in Appendix).
@@ 3.2 Cookbook
Our model, Kyvo, starts from language as the principal modality. Our backbone is the decoder-only Llama-3.2-1B-Instruct [grattafiori2024llama3herdmodels] initialized from language-only pretrained weights. We extend it with modality-specific tokenizers for images and the structured 3D modality, along with modified input embeddings and output projections (Fig.˜1). We do this in order to leverage the generalization and reasoning capabilities of LLMs, a hypothesis we later confirm in Tab.˜5(b). We evaluate key architectural choices – each with significant performance impact – which we hope offer insights to guide future multimodal LLM development in the 3D domain.
We begin by exploring how to represent and tokenize our structured 3D modality. We explore the tokenization of individual 3D objects of diverse geometry and appearance from Objaverse as well as the tokenization of scenes containing multiple objects of known geometry in CLEVR. This setup helps establish best practices for integrating 3D into a unified LLM framework. We then extend these findings to complex scenes in ObjaWorld (Sec.˜3.3 & Sec.˜3.4) and validate generalization to real-world scenes in Objectron and ARKitScenes (Sec.˜3.5).
3.2.1 What is the optimal 3D representation?
Our model handles three modalities – images, text, and structured 3D scenes – by converting each into token sequences for autoregressive modeling.
Text. We employ an off-the-shelf text tokenizer from Llama-3.2 [grattafiori2024llama3herdmodels] with a 128,000128,\!000 size vocabulary. Text instructions, questions, and answers are tokenized and enclosed within special, learnable tokens [TEXT-START] and [TEXT-END].
Images. Our framework addresses tasks involving images as both inputs and outputs. To enable image generation within an autoregressive paradigm, we adopt discrete image representations using VQGAN [esser2021taming]. This approach maps continuous image features to discrete tokens through a learned codebook. Specifically, we fine-tune a pre-trained VQGAN model to optimize the codebook for our visual domain. Image tokens are enclosed within special, learnable tokens, [IMAGE-START] and [IMAGE-END].
3D scene tokenization. We represent 3D scenes as structured token sequences, where each list element encodes one object. Object attributes – shape, size, location, color, material – are expressed through special learnable tokens (e.g., [SHAPE]), whose values may be text (“car”, “yellow”), numbers (for size or position), or learned 3D embeddings. Each object is enclosed by [OBJECT-START] and [OBJECT-END], and the full scene by [SCENE-START] and [SCENE-END]. An example scene with two objects is shown below:
@@ 3.2.1 What is the optimal 3D representation?
We discuss two key design decisions:
Figure 3: 3D VQ-VAE training involves the standard VQ-VAE losses (including reconstruction loss) applied in latent space as well as an auxiliary reconstruction loss applied in decoded pixel space.
[…]
Figure 4: 3D Tokenization findings and comparisons. (a) Effect of auxiliary reconstruction loss. An auxiliary pixel-space reconstruction loss on decoded renders from multiple views of the 3D object leads to much better reconstructions. (b) 3D tokenizer comparison. Reconstructions from our Trellis-based VQ-VAE exceed the quality of SAR3D reconstructions with fewer tokens; improved textures stem from the Trellis slat representation rather than triplanes. (c) Learned 3D shape encodings are effective during decoding. The 3D tokens used in Kyvo are sufficient for both reconstruction and rendering using Llama 3.2 as decoder.
Aux Rec Loss Mean Rank()($\downarrow$})
$2.828$
Fixed Single View $1.672$
Multiple Views $\mathbf{1.500}$
Table 1: Effect of auxiliary reconstruction loss.
3D Tokenizer Mean Rank()($\downarrow$})
SAR3D $1.605$
Kyvo 3D VQ-VAE $\mathbf{1.395}$
Table 2: 3D tokenization comparison using human evaluation.
(1) How to tokenize 3D shapes (geometry & texture)? We aim for our structured 3D modality to encode complex objects via compact, autoregressively decodable 3D shape representations. To this end, we adopt Trellis [xiang2024structured], which encodes geometry and texture as sparse voxel features (slats) $z=\{(z_{i},p_{i})\}_{i=1}^{L}$, where each $z_{i}\in\mathbb{R}^{8}$ represents local features and $p_{i}$ indexes active voxels in an $N^{3}$ grid. Although slats are sparse ($L\ll N^{3}$, typically $L\approx 20\text{k}$ for $N=64$), their length makes autoregressive modeling intractable.
We therefore train a 3D VQ-VAE [van2017neural] to compress slats from $64^{3}\times 8$ to a dense $8^{3}\times 128$ latent, vector-quantized with an 8192-token codebook. Each object is thus represented by 512 discrete tokens—a $\sim\!40\times$ reduction—enabling efficient autoregressive decoding.
How do we train such a 3D VQ-VAE in latent slat space while preserving essential geometric and appearance information? We find that training the VQ-VAE in the latent space of slats is insufficient to learn an effective representation for reconstruction. To overcome this, we apply an auxiliary reconstruction loss to the decoded reconstructed slats in pixel-space, as shown in Fig.˜3. We use the same pixel-space reconstruction loss ($\mathcal{L}_{1}$, D-SSIM and LPIPS) as in [xiang2024structured]. This yields significantly better reconstructions, despite achieving a similar latent space reconstruction loss, as shown in Fig.˜4. Additionally, we find that imposing a multi-view reconstruction loss of the asset (randomly sampled from 150 views) leads to better reconstructions than a single fixed view. We quantify the improvement using human evaluations in Tab.˜2.
Our Trellis-based 3D VQ-VAE matches or surpasses SAR3D [chen2024sar3d] while using 4× fewer tokens (512 vs. 2040). Qualitative and quantitive comparisons are shown in Fig.˜4 and Tab.˜2. Importantly, our compact 3D tokens integrate seamlessly into our unified vocabulary with image and language tokens, enabling efficient multi-object scene encoding and autoregressive decoding, while SAR3D tackles single-asset generation only.
While our representation enables effective 3D asset tokenization and reconstruction, we further test its suitability for autoregressive decoding. Using the same Llama-3.2-Instruct backbone, we evaluate reconstruction and rendering on unseen assets, in Fig.˜4. The learned 3D encodings generalize well, confirming their effectiveness for both reconstruction and decoding. These 3D tokens are then incorporated into our unified vocabulary alongside image and language tokens for training and inference.
(2) How to tokenize 3D location and orientation? In addition to 3D shape, the 3D location and orientation of an object are key attributes within our 3D modality and critical for understanding 3D spatial relationships and performing grounded 3D actions based on instructions. Thus, accurately encoding coordinates is essential. However, LLMs are struggle with numbers [mirzadeh2024gsm, singh2024tokenization].
Granularity Rendering()($\downarrow$}) Recognition()($\uparrow$}) Instruction()($\uparrow$}) QA()($\uparrow$})
@@
To mitigate this, we encode each object’s $x,y,z$ coordinates separately as individual tokens, allowing the model to learn distinct embeddings for each coordinate. We discretize the coordinate values using equally spaced bins based on a chosen granularity. We find this granularity to be decisive for performance – if it is too coarse, spatial inaccuracies arise; if it is too fine, it leads to an exponential increase in tokens, with fewer training samples per bin and thus difficulty in learning. Tab.˜3 shows the effect of granularity across the four tasks on CLEVR. We choose CLEVR, which features simple, known shapes, to isolate and study the effect of number encodings in next-token prediction frameworks. We find that a granularity of $0.05$ outperforms the coarser $0.5$ and the finer $0.005$.  Fig.˜5 compares image generation for the rendering task on the test set across varying levels of granularity.
@@
Furthermore, our tokenization approach substantially improves efficiency by reducing sequence length compared to standard text tokenizers, which typically fragment floating-point values (e.g., "0.0000.000" becomes "0", ".", "000000"). We achieve a mean sequence length reduction from $271.4$ tokens using the standard Llama-3.2 tokenizer to $93.2$ tokens – a $2.91\times$ compression ratio. These efficiency gains directly translate to decreased memory requirements and faster training and inference times. The final vocabulary comprises 137,607137,\!607 tokens covering all modalities.